# Using Diffusion Models to do Data Assimilation

### v1

@@ 4 Numerical illustration
We illustrate the various forms of Bayes’ rule and their corresponding diffusion models of Section 3 with a specific example.
@@ 4.1 Discretization of the dynamical model
4.2 Data assimilation
4.2.1 Kalman filtering
The Kalman filter with a climatological prior evaluates the equation for the posterior mean (23) at each cycle. The Kalman filter with a cycling prior (Section 3.3.2.3.2.1) is initialized with the climatological moments of $\mathbf{P}_{0}=\mathbf{P}_{c}$ and $\overline{\mathbf{x}}_{0}=\mathbf{0}$. The discrete model is used to propagate the posterior mean and covariance (see equations (75) and (76)) forward in time to the next set of observations. The Kalman filter with an extended likelihood uses the posterior mean and covariance equations in Section Section 3.3.3.3.3.1. Moreover, at each cycle we use the discretized model (equation (73)) to integrate the posterior mean forward in time to the next cycle.
Finally, we must specify the parameters $a$ and $r_{f}$ for the Kalman filter with an extended likelihood. We leverage the following strategy with the ultimate goal of finding a combination of $a$ and $r_{f}$ such that the posterior mean squared error (MSE) of the Kalman filter matches its posterior variance prediction, which is key to verifying that the theory is matching the experiment. We initialize the process with $a=1$ and $r_{f}=1$ and cycle the corresponding Kalman filter for $10^{5}$ times to collect truth and forecast pairs $([\mathbf{x}_{k}]^{j},[\mathbf{f}_{k}]^{j})$ at each grid point $j$. We then fit two functions to this large dataset of forecast-truth pairs. First, we can fit a line through the data because the mean of $p(\mathbf{f}_{k}|\mathbf{x}_{k})$ is $a\mathbf{x}_{k}$, so that the slope of the line can be used as a candidate value for $a$. Second, we can compute the MSE associated with the forecast, i.e.,
@@ 4.2.1 Kalman filtering
and subsequently fit a quadratic to obtain another estimate of the parameter $a$ and also an estimate of the parameter $r_{f}$. We then repeat this process by re-running the cycling experiment with the new values of $a$ and $r_{f}$ until the two estimates of $a$ agree to two decimal places. As an example, for the case of $\Delta=0.1$, the system settled on $a=0.61$ and $r_{f}=0.34$ and the time-averaged MSE of the Kalman filter was then very close to the posterior variance prediction. We realize that this strategy is likely to not be practical in typical geophysical systems. The goal here is as a benchmark to the associated diffusion model and not as a practical algorithm.
4.2.2 Diffusion models
We build a diffusion model using a cycling prior of the form (49) using a simple FE-based method, viz.
$\mathbf{v}_{n-1}=\mathbf{v}_{n}-\frac{2}{t_{n}}\Big{[}\mathbf{P}_{i}^{f}\hat{\mathbf{H}}^{T}[\hat{\mathbf{H}}\mathbf{P}_{i}^{f}\hat{\mathbf{H}}^{T}+\mathbf{R}]^{-1}\hat{\mathbf{y}}-(\mathbf{v}_{n}-\overline{\mathbf{x}}_{i})\Big{]}\Delta_{d}+\sqrt{2t_{n}|\Delta_{d}|}\mathbf{w}_{n}$ (79)
@@ 4.2.2 Diffusion models
$\hat{\mathbf{y}}=\begin{bmatrix}\mathbf{v}-\overline{\mathbf{x}}_{i}\\ \mathbf{y}_{k}-\mathbf{H}\overline{\mathbf{x}}_{i}\end{bmatrix}$ (80)
with $\hat{\mathbf{H}}$ and $\mathbf{R}$ being defined as in section 3.2.2 and we generate the $i^{th}$ sample as $\mathbf{x}^{(i)}=\mathbf{v}_{0}$. We use 1000 internal virtual time steps denoted by the subscript, $n$, for the reverse integration with a time step of $\Delta_{d}=-0.1$ and we repeat this to create an ensemble of $n_{e}=10^{4}$ members. Because the reverse process must start from a large virtual time we integrate this equation starting at an arbitrarily chosen time of $T=100$ back to $t=0$ on a uniform temporal grid. From these ensemble members we then calculate a posterior ensemble mean and covariance matrix. Lastly, we then propagate this posterior ensemble mean and covariance matrix forward to the next set of observations using (75) and (76).
For the climatological diffusion model we also solve (31) using a simple FE-based method, viz.
$\mathbf{v}_{n-1}=\mathbf{v}_{n}-\frac{2}{t_{n}}\Big{[}\mathbf{P}_{c}^{f}\hat{\mathbf{H}}^{T}[\hat{\mathbf{H}}\mathbf{P}_{c}^{f}\hat{\mathbf{H}}^{T}+\mathbf{R}]^{-1}\hat{\mathbf{y}}-\mathbf{v}_{n}\Big{]}\Delta_{d}+\sqrt{2t_{n}|\Delta_{d}|}\mathbf{w}_{n}$ (81)
where
$\hat{\mathbf{y}}=\begin{bmatrix}\mathbf{v}\\ \mathbf{y}_{k}\end{bmatrix}$ (82)
with $\hat{\mathbf{H}}$ and $\mathbf{R}$ being defined as in section 3.1.2. We again create $n_{e}=10^{4}$ members from which we determine the posterior ensemble mean and covariance that we report below. In distinction to the diffusion model with the cycling prior these posterior ensemble mean and covariances are not used in any way at the next cycle.
For the extended likelihood diffusion model we solve (68) using a FE method as
$\mathbf{v}_{n-1}=\mathbf{v}_{n}-\frac{2}{t_{n}}\Big{[}\mathbf{P}_{c}^{f}\hat{\mathbf{H}}^{T}[\hat{\mathbf{H}}\mathbf{P}_{c}^{f}\hat{\mathbf{H}}^{T}+\mathbf{R}]^{-1}\hat{\mathbf{y}}-\mathbf{v}_{n}\Big{]}\Delta_{d}+\sqrt{2t_{n}|\Delta_{d}|}\mathbf{w}_{n}$ (83)
where
$\hat{\mathbf{y}}=\begin{bmatrix}\mathbf{v}\\ \mathbf{y}_{k}\\ \mathbf{f}_{k}\end{bmatrix}$ (84)
with $\hat{\mathbf{H}}$ and $\mathbf{R}$ being defined as in section 3.3.2. These matrices require knowledge of $a$ and $r_{f}$, for which we use the same values determined using the Kalman filter setup. Similar to the diffusion model using the climatological prior we create $n_{e}=10^{4}$ members from which we determine the posterior ensemble mean and covariance. However, in distinction to the diffusion model using the climatological prior we now take the posterior mean and integrate it forward in time to the next set of observations using (73) to obtain the forecast, $\mathbf{f}_{k}$.
4.2.3 Results
We first run an experiment with $\Delta=0.1$ in which we observe every state element ($\mathbf{H}=\mathbf{I}$) and collect observations at every time step of the model; the observation error variance is $r=1$. This experiment, and the others reported below, will use a set of $n_{y}=100$ observations. The posterior MSE and posterior variance are summarized in Figure 1, for three DA setups (climatological prior, extended likelihood and cycling prior), solved analytically via the corresponding Kalman formalisms, or via a diffusion model.
Figure 1: (a) MSE (solid or dashed) and variance (dotted) for three different Bayesian posterior distributions, corresponding to a climatological prior, a cycling prior and an extended likelihood. Shown are results obtained with Kalman filters (analytical solutions, solid) and diffusion models (dashed). Note that the MSE curves for the Kalman filter methods coincide with those of the diffusion models, indicating that the two methods produce identical results. The posterior variances are slightly different between the Kalman and diffusion-based methods because of small errors due to the choice of the FE method for the solution of the diffusion models’ SDEs. (b)-(d) Histograms of the MSE after a 20-cycle spin-up period. Again, the diffusion models produce nearly identical results as the Kalman filter methods, and we see that MSE decreases (on average and in distribution) when moving from a DA system with a climatological prior, to one with an extended likelihood, to one with a cycling prior.
As expected from our theory, we observe that (i) the diffusion models successfully emulate the corresponding Kalman methods, i.e., the diffusion models emulate their associated form of Bayes’ rule; (ii) a DA system with a cycling prior generates the smallest MSE and posterior variance, while DA systems with a climatological prior or an extended likelihood lead to larger errors; (iii) bringing in a forecast via an extended likelihood improves the state estimates when compared to a DA system with a climatological prior, but the errors and variances are still larger than what a fully cycled DA system can achieve.
This example provides a numerical confirmation of our theory. In particular, this example shows that the various forms of Bayes’ rule being described by the Kalman filters are in fact realized by the particular choices used to construct the training sets used to create the three diffusion models. The posterior using climatological prior realized via a Kalman filter or diffusion model produce the same moments for the posterior and, hence, the two systems are indeed working with the same form of Bayes’ rule. Analogous results hold for the posteriors defined by a cycling prior or an extended likelihood. It is also important to check that the posterior variance correctly matches the time-averaged MSE for all methods, which is the most important indicator that the theory is working correctly.
Intuitively, as the time interval between observations, $\Delta$, becomes larger, past observations carry less information and thus one would expect that the cycling prior converges to the climatological prior as $\Delta$ becomes large. We now test this and other notions in numerical experiments in which we vary $\Delta$ and the observation error variance, $r$, and then compare the steady-state posterior covariance of a DA system with a cycling prior to that of a DA system with a climatological prior.
Figure 2: (a) Steady state posterior variance of a DA system with a cycling prior as a function of observation error and time interval between observations. Also shown is the steady state posterior variance of a DA system with a climatological prior, which corresponds to a very large timer interval between observations. (b) Steady state posterior variance of a DA system with an extended likelihood as a function of observation error (time interval between observations is $\Delta=0.1$). Also shown are the posterior variances of a DA system with a cycling prior and with a climatological prior.
To this end, we repeat these cycling DA experiments for various observation error variances ($r$) and summarize our results in Figure 2(a). Note that as the time between observations increases from $\Delta=0.01$ to $\Delta=2$, the posterior variance using the cycling prior indeed converges to the posterior variance using the climatological prior. Hence, the time between observations strongly controls the difference between those two forms of Bayes’ rule and the resulting DA systems. More specifically, there is a delicate balance between the timescale of the error growth induced by the stochastic term in (17) and the variance reducing property of the assimilation of observations. If the error growth dominates (large time interval between observations), DA systems with a cycling prior are nearly identical to DA systems with a climatological prior. If the assimilation of observations is frequent (short time interval between observations), then the cycling prior propagates information from past DA cycles to the current ones and, therefore, reduces state estimation errors and posterior variances.
Finally, we perform a set of experiments in which we vary the observation error variance for the DA system with an extended likelihood. We keep the time interval between observations fixed (at $\Delta=0.1$) because for each value of $r$ we need to tune the parameters $a$ and $r_{f}$ of the extended likelihood system. Results are summarized in Figure 2(b), where we show the posterior variance of an extended likelihood DA system as a function of the observation error variance. For comparison, we also plot the posterior variance of a DA system with a climatological prior and with a cycling prior (already shown in Figure 2(b)).
We note that the posterior variance of a DA system with an extended likelihood is in between that of a system with a cycling or climatological prior, unless $r$ is very large (see below). Thus, for moderate $r$, the extended likelihood indeed propagates information from previous assimilation steps via a single forecast. The posterior variance of the extended likelihood DA system, however, is larger than that of a DA system with a cycling prior, which indicates that more information is transmitted between cycles when the entire distribution is propagated, rather than a single forecast. Moreover, the posterior variance of a DA system with an extended likelihood converges to the posterior variance of a DA system with a climatological prior as the observation error variance $r$ becomes large. This is due to the fact that the information from the forecast is only as good as the information in the posterior mean it is integrated from. As the observation error increases we find that $r_{f}$ also increases such that there is less information in the forecast and therefore the differences between the posterior using a climatological prior and the posterior using the extended likelihood become muted. Note, however, that there is still a large difference between the posterior variance using a cycling prior and the posterior variance using the climatological prior even for very large observation error variances. Hence, we again see that propagating the entire distribution leads to more accurate state estimates than propagating a single forecast.
5 Summary and conclusions
We have shown that a diffusion DA system can target different Bayesian posterior distributions and we have explored, in detail, three versions: a Bayesian posterior with a cycled prior, which is typical of conventional DA systems (equation (1)), a Bayesian posterior with a climatological prior (equation (14)), and a Bayesian posterior in which the likelihood is extended with a single forecast (equation (58)) that is treated like an additional predictor. We have shown three different ways to construct training sets that lead to diffusion models that sample each of these three Bayesian posterior distributions.
The key aspect of the differences between these versions of Bayes’ rule is in their use of the prior. In both the posterior using a climatological prior and the posterior using an extended likelihood the training set is determined from a long time-series of past weather. This long time-series of past weather constitutes random samples from a climatological prior. This climatological prior is never updated in these versions of the posterior. Hence, both these systems ostensibly require only a single training. However, we envision two possible ways one might make use of the posterior using an extended likelihood. In the first way, one might run the extended likelihood diffusion model alongside a conventional DA system. This conventional DA system would produce the forecast used in the extended likelihood diffusion model. Because the quality of the forecast from the conventional DA system is stationary the training set will correctly provide the appropriate examples. However, the second way one might make use of the posterior using an extended likelihood is with a self-sufficient diffusion model that produces its own forecasts. We speculate that the iterative procedure we used to find $a$ and $r_{f}$ implies that using the posterior with an extended likelihood in a diffusion model will require multiple training attempts to get the forecasts produced from the diffusion-based DA system to be of the same quality as the ones trained upon. Therefore, in this second case the posterior using the extended likelihood is likely to be more computationally demanding in order to obtain a near optimal system. Of course, one could ignore this warning and simply train this extended likelihood diffusion model once, but in our experiments (not shown) we found this to lead to a system that produced ensembles whose variance was not as carefully calibrated to the MSE of the ensemble mean. Furthermore, the posterior using a cycled prior quite obviously updates its prior training set at each and every cycle. This requires re-training of the diffusion model at every cycle, and is therefore the most computationally demanding system we examined. Given the number of samples required to train typical diffusion model architectures this implies a significant expense in both the generation of the new training set as well as in the training itself.
Nevertheless, the posterior using a cycled prior had nearly universally lower errors than the other two posterior distributions. Only when the observations were so far apart in time that the observations entirely de-correlated did all three versions of Bayes’ rule deliver a similar answer. Hence, we suggest that future research should be directed towards ways in which we can make use of the posterior with a cycling prior in the most efficient ways possible. For example, it may be that the difference between the diffusion model at cycle $k-1$ and $k$ is small enough that one could make use of fine-tuning methods (see e.g. Parthasarathy et al. (2024)) and/or transfer learning (see e.g. Zhuang et al. (2020)). Similarly, future research into variations on the likelihood based approximations in Chung et al. (2023) as applied to the posteriors for the extended likelihood and cycling prior may also lead to ways in which we can accelerate the training required. Work in these directions is already underway.
Acknowledgments
DH is supported by the US Office of Naval Research (ONR) grant N0001422WX00451. MM is supported by the US ONR grant N000142512298.
Data Statement
The code used for making the figures will be made available on github.
@@ Acknowledgments
Appendix A Tweedie’s formula
For Tweedie’s formula, we consider a random variable $v$, conditioned on a random variable $x$, such that
$v|x=x+\varepsilon,\quad\varepsilon\sim\mathcal{N}(0,\sigma_{0}^{2}),$ (85)
which implies the conditional pdf
$p(v|x)=\frac{1}{\sqrt{2\pi}\sigma_{0}}\exp\left(-\frac{1}{2}\left(\frac{v-x}{\sigma_{0}}\right)^{2}\right).$ (86)
We wish to compute
$\frac{\text{d}}{\text{d}v}\log p(v)=\frac{1}{p(v)}\frac{\text{d}}{\text{d}v}p(v).$ (87)
Using the fact that
$p(v)=\int_{-\infty}^{\infty}p(v|x)p(x)\text{d}x,$ (88)
we find
$\frac{\text{d}}{\text{d}v}p(v)=\int_{-\infty}^{\infty}\frac{x-v}{\sigma_{0}^{2}}p(v|x)p(x)\text{d}x.$ (89)
Thus
$\frac{\text{d}}{\text{d}v}\log p(v)=\int_{-\infty}^{\infty}\frac{x-v}{\sigma_{0}^{2}}\frac{p(v|x)p(x)}{p(v)}\text{d}x,$ (90)
which simplifies to
$\frac{\text{d}}{\text{d}v}\log p(v)=\int_{-\infty}^{\infty}\frac{x-v}{\sigma_{0}^{2}}p(x|v)\text{d}x,$ (91)
since
$\frac{p(v|x)p(x)}{p(v)}=\frac{\frac{p(v,x)}{p(x)}p(x)}{p(v)}=\frac{p(v,x)}{p(v)}=p(x|v).$ (92)
Distributing the integral leads to
$\frac{\text{d}}{\text{d}v}\log p(v)=\frac{1}{\sigma_{0}^{2}}\int_{-\infty}^{\infty}xp(x|v)\text{d}x-v\int_{-\infty}^{\infty}p(x|v)\text{d}x=\frac{1}{\sigma_{0}^{2}}\left(E[x|v]-v\right),$ (93)
which, upon rearrangement, gives Tweedie’s formula:
$E[x|v]=v+\sigma_{0}^{2}\frac{\text{d}}{\text{d}v}\log(p(v)).$ (94)
Appendix B Solution to the SDE in (32)
We begin with (32) and perform a few simple manipulations:
@@ Appendix B Solution to the SDE in (32)
Integrating backwards in time, from $T$ to 00,
@@
which is the desired result.
Appendix C Solution to the SDE in (50)
To solve the SDE
@@ Appendix C Solution to the SDE in (50)
Integrating backwards in time, from $T$ to 00,

### v2

@@ 5 Numerical illustration
We illustrate the variants of diffusion DA systems with a specific example, which also re-iterates the connections with variants of Bayesian posterior distributions. In the diffusion modeling literature, forward and reverse processes are derived in continuous time and subsequently discretized. We follow the same approach and discretize forward and reverse processes using simple forward Euler methods. We discretize the model (23) with a similar forward Euler scheme, but other methods, or even analytical solution, will yield similar results.
@@ 5.1 Discretization of the dynamical model
5.2 Diffusion DA implementation details
We implement all three diffusion DA systems and discretize their reverse processes using a forward Euler scheme to integrate from $T=100$ to $t=0$; the time step is set to $0.1$. At each cycle, we generate an ensemble of $10^{4}$ samples from which we compute the mean and variances. Note that the diffusion DA system with a cycling prior re-uses the ensemble from the previous cycle for re-training. The diffusion DA with an extended likelihood uses the posterior ensemble mean from the previous cycle to generate a single forecast, i.e. it does not propagate an ensemble forward in time.
The diffusion DA with an extended likelihood further requires that we specify the extended likelihood parameters $a$ and $r_{f}$. To find these, we use Kalman filtering as follows. We initialize the process with $a=1$ and $r_{f}=1$ and cycle the Kalman filter with extended likelihood (Section 4.4.4.4.4.1) for $10^{5}$ cycles to collect truth and forecast pairs $([\mathbf{x}_{k}]^{j},[\mathbf{f}_{k}]^{j})$ at each grid point $j$. We then fit two functions to this large dataset of forecast-truth pairs. First, we can fit a line through the data because the mean of $p(\mathbf{f}_{k}|\mathbf{x}_{k})$ is $a\mathbf{x}_{k}$, so that the slope of the line can be used as a candidate value for $a$. Second, we can compute the MSE associated with the forecast, i.e.,
@@ 5.2 Diffusion DA implementation details
and subsequently fit a quadratic to obtain another estimate of the parameter $a$ and also an estimate of the parameter $r_{f}$. We then repeat this process by re-running the cycling experiment with the new values of $a$ and $r_{f}$ until the two estimates of $a$ agree to two decimal places. As an example, for the case of $\Delta=0.1$ time units between observations, the system settled on $a=0.61$ and $r_{f}=0.34$ and the time-averaged MSE of the Kalman filter was then very close to the posterior variance prediction. Lastly, we emphasize that this complicated iterative process for obtaining these two parameters is only necessary here because we avoided neural networks for the denoiser. If we had used a neural network, rather than analytic formulas, a sufficiently expressive model would have learned these parameters from the data.
5.2.1 Results
We first run one experiment to confirm our theory. Here, we observe every state element ($\mathbf{H}=\mathbf{I}$) every $\Delta=0.1$ time units and choose the observation error variance $r=1$. The posterior MSE and posterior variances are summarized in Figure 1, for the three diffusion DA systems (climatological prior, extended likelihood and cycling prior).
Figure 1: MSE (thick) and variance (thin) for diffusion DA systems with a climatological prior (green), an extended likelihood (orange) and a cycling prior (purple). MSE and variances from the Bayesian solutions nearly coincide with those of the diffusion DA systems, indicating that the two methods produce nearly identical results.
As expected from our theory, we observe that (i) a diffusion DA system with a cycling prior generates the smallest MSE and posterior variance, while DA systems with a climatological prior or an extended likelihood lead to larger errors; (ii) bringing in a forecast via an extended likelihood improves the state estimates when compared to a DA system with only a climatological prior, but the errors and variances are still larger than what a fully cycled DA system can achieve; (iii) posterior variances match time-averaged MSE for all three diffusion DA systems, which is an important indicator that our theory is working correctly. Thus, our example provides a numerical confirmation of our theory that the diffusion models with three different training datasets indeed realize three different Bayesian posterior distributions.
We perform additional numerical experiments in which we vary the time between observations $\Delta$ or the observation error $r$. Intuitively, as the time interval between observations, $\Delta$, becomes larger, past observations carry less information and thus one would expect that the cycling prior converges to the climatological prior as $\Delta$ becomes large. We test this idea for various choices of $r$, with results summarized in Figure 2(a).
Figure 2: (a) Steady state posterior variance of a DA system with a cycling prior as a function of observation error and time interval between observations. Also shown is the steady state posterior variance of a DA system with a climatological prior, which corresponds to a very large time interval between observations. (b) Steady state posterior variance of a DA system with an extended likelihood as a function of observation error (time interval between observations is $\Delta=0.1$). Also shown are the posterior variances of a DA system with a cycling prior and with a climatological prior.
As the time between observations increases from $\Delta=0.01$ to $\Delta=2$, the posterior variance corresponding to a cycling prior converges to the posterior variance corresponding to a climatological prior (independently of $r$). Hence, the time between observations strongly controls the difference between those two forms of Bayesian posterior distributions and the resulting DA systems. More specifically, there is a delicate balance between the timescale of the error growth induced by the stochastic term in (23) and the variance reducing property of the assimilation of observations. If the error growth dominates (large time interval between observations), DA systems with a cycling prior are nearly identical to DA systems with a climatological prior. If the assimilation of observations is frequent (short time interval between observations), then the cycling prior propagates information from past DA cycles to the current ones and, therefore, reduces state estimation errors and posterior variances.
Next, we perform a set of experiments in which we vary the observation error variance for a DA system with an extended likelihood. Results for $\Delta=0.1$ are summarized in Figure 2(b), where we show the posterior variance of an extended likelihood DA system as a function of the observation error variance. For comparison, we also plot the posterior variance of a DA system with a climatological prior and with a cycling prior (already shown in Figure 2(a)). The posterior variance of a DA system with an extended likelihood is in between that of a system with a cycling or climatological prior, unless $r$ is very large (see below). Thus, for moderate $r$, the extended likelihood indeed propagates information from previous assimilation steps via a single forecast. The posterior variance of the extended likelihood DA system, however, is larger than that of a DA system with a cycling prior, which indicates that more information is transmitted between cycles when the entire distribution is propagated via an ensemble (i.e. a time-evolving training set), rather than a single forecast. Moreover, the posterior variance of a DA system with an extended likelihood converges to the posterior variance of a DA system with a climatological prior as the observation error variance $r$ becomes large. This is due to the fact that the information from the forecast is only as good as the information in the posterior mean it is integrated from. As the observation error increases we find that $r_{f}$ also increases such that there is less information in the forecast and therefore the differences between the posterior using a climatological prior and the posterior using the extended likelihood become muted. Note, however, that there is still a large difference between the posterior variance using a cycling prior and the posterior variance using the climatological prior even for very large observation error variances. Hence, we again see that propagating the entire distribution leads to more accurate state estimates than propagating a single forecast.
Figure 3: (a) Gain (dashed) and posterior variance (solid) for a diffusion DA system using a climatological prior as a function of observation error variance. Shown are the gain/variance of diffusion DA systems with climatological prior and an exact denoiser (green) or approximate denoiser (black, see Section 4.b.3). (b) MSE (thick) and variance (thin) for diffusion DA systems with a climatological prior (green) in a cycling DA experiment similar to Figure 1. Here we compare the climatological diffusion DA system using the exact denoiser (green) to the approximate denoiser (black).
Lastly, we compare diffusion DA systems with a climatological prior using the exact and approximate denoisers (Section 4.b.3). In Figure 3(a) we show the gain and posterior variance as a function of the observation error variance ($r$). Here, the “gain” is the quantity multiplying the observation when we compute the mean (compare to (31)), i.e., for the system with the exact denoiser, the gain is given by (45) and for the approximate denoiser the gain is (56). The variance of the system with the exact denoiser is given by (49) and the variance of the system with the approximate denoiser is (59). From this plot one can see that the gain of the system with the approximate denoiser is biased too large and therefore draws too closely to the observations. This effect is strongest when the observation error is much smaller than the climatological error variance. Similarly, the posterior variance produced by the approximate denoiser is too small. These properties of the approximate denoiser transfer directly to explaining its behavior when used in a cycling DA system as shown in Figure 3(b). Here, we show MSE and variance as a function of DA cycle for diffusion DA systems with approximate and exact denoisers (both using climatological priors). We see that the approximate denoiser has excess MSE as well as a posterior variance that is too small. Both of these properties are consistent with the biased gain and biased posterior variance.
6 Similarities and differences between diffusion DA, ensemble DA and variational DA
Our analysis and numerical experiments reveal the relationship between the samples one receives from a particular diffusion DA system and the prior distribution used in training. When one uses a climatological prior for training then that diffusion DA system is trained only once. Additionally, we can use an approximate denoiser to avoid re-training when applying this DA system to new observational platforms, but the approximation in the denoiser results in unavoidable biases. A diffusion DA system whose denoiser is retrained at every cycle samples a Bayesian posterior distribution with a cycling prior (the commonly used prior in ensemble DA). A diffusion DA system can be modified to ingest a forecast along with the observations. Such a system samples a Bayesian posterior distribution somewhere in between the one with a climatological and cycling prior.
Since diffusion DA and ensemble DA are sampling algorithms, connections between such systems are rather obvious. In particular, a diffusion DA system whose denoiser is re-trained at every cycle is equivalent to a traditional ensemble DA system, but it may be more robust to nonlinearity or non-Gaussianity. There is no traditional, equivalent ensemble DA system that replicates the diffusion DA system that samples a Bayesian posterior distribution with a climatological prior. The reason is likely that such a system is suboptimal, as our numerical examples show, because a cycling DA system achieves a smaller error and smaller variance due to the cycling prior. From an ensemble DA perspective, diffusion DA systems trained on a long dataset (climatological prior), are thus a rather unusual idea, but nevertheless these systems do seem to be a natural first choice for ML/AI modeling. With that said, our analysis and numerical experiments suggest that such systems may not be able to achieve the accuracy of a DA system that uses a cycled prior. We find the fact that diffusion DA systems with a climatological prior are theoretically sub-optimal, yet producing comparable results to cycled DA in practice, very interesting. One reason could be that the prior distributions found in NWP may have substantial non-Gaussianity, and diffusion DA methods may be able to account for non-Gaussianity more accurately than traditional ensemble DA.
Diffusion models are sampling tools, while variational methods rely on optimization and only produce a single, optimal state estimate. Therefore, the connections between variational DA and diffusion modeling are somewhat murky, yet many papers on diffusion DA invoke similarities with 3D-Var or 4D-Var (e.g., Manshausen et al. (2024)). The focus of our paper is diffusion DA and the various Bayesian posterior distributions one can sample with these techniques and corresponding training datasets. However, we can also comment on similarities and differences between diffusion DA and variational DA. One key difference here is how the background covariance matrix is constructed. Traditional variational DA uses a constant-in-time background covariance. But this background covariance is not the climatological covariance, but rather a covariance designed to represent short-term forecast errors (Parrish and Derber (1992)). If one were to try to make a loose connection between variational and diffusion DA using a climatological prior, then one could argue that both methods use a static, time-independent prior or, equivalently, training dataset. The training dataset used in diffusion DA (the climatological prior), however, does not correspond to the prior typically used in variational DA. Instead, the long training dataset used in this form of diffusion DA corresponds to an unusual and, as we have shown, suboptimal choice of the prior.
7 Summary and conclusions
We have considered the question of how to use diffusion modeling to do data assimilation. Specifically, we have worked out how different training datasets in diffusion modeling correspond to different prior assumptions and, hence, how these different diffusion DA systems sample from different Bayesian posterior distributions. Our main results can be summarized as follows.
  1. 1.
    Diffusion DA is effective at sampling a Bayesian posterior distribution with a fixed, climatological prior, but the accuracy of such a system is inferior to an ensemble DA system that samples a Bayesian posterior distribution with a time-evolving, cycling prior that propagates information from one DA cycle to the next. Indeed, the training dataset in diffusion DA corresponds to a prior that is rather unusual in traditional DA.
  2. 2.
    A diffusion DA system can sample the same posterior distribution as an ensemble DA system using a time-evolving cycling prior, but the denoiser in such a diffusion DA system will need to be re-trained at each DA cycle.
  3. 3.
    A diffusion DA system with a fixed, climatological prior can be modified to ingest a forecast in addition to the observations. This forecast adds some aspects of the time-evolving cycling prior back into the DA system, but is not entirely equivalent to a fully cycling ensemble DA. If the forecast is generated by a separate DA system, the training cost is relatively low and the accuracy is in between the accuracy of a DA system with a climatological prior and a DA system with a cycling prior.
  4. 4.
    One can make use of an approximate denoiser when using the climatological prior. While this results in a diffusion DA system that does not need re-training when making use of a new observation network, it does incur a penalty of a biased gain and a biased posterior variance. Additionally, while it would be difficult to alleviate the bias in the gain, the posterior variance issue is likely to be easily fixed by the commonly available tunable parameters in the reverse diffusion process. Lastly, because one already re-trains the diffusion model at every cycle when using the time-evolving, cycling prior the computational benefits of the approximate denoiser do not seem to apply.
Our results suggest that efficient implementation of diffusion DA systems with a cycling prior may yield better results than current diffusion DA systems. Possibilities for efficient training include fine-tuning methods (see, e.g., Parthasarathy et al., 2024)), transfer learning (see, e.g., Zhuang et al., 2020)), or likelihood based approximations (see, e.g., Chung et al., 2023). Work in these directions is already underway.
Acknowledgements.
[…]
$z=u|y+\epsilon$ (97)
where $\epsilon$ is a draw from $N(0,\sigma_{o}^{2})$, which is assumed independent of $y$. This implies that we can write
$p(z|y)=\int_{-\infty}^{\infty}p(z|u)p(u|y)du$ (98)
@@ Acknowledgements.
$p(z|u)=\frac{1}{\sqrt{2\pi\sigma_{o}^{2}}}exp\Big{(}-\frac{1}{2\sigma_{o}^{2}}(z-u)^{2}\Big{)}$ (99)
Similarly, we can find the derivative as
$\begin{split}\frac{d}{dz}(log(p(z|y)))=\frac{1}{p(z|y)}\frac{dp(z|y)}{dz}\\ =-\int_{-\infty}^{\infty}\frac{(z-u)}{\sigma_{o}^{2}}\frac{p(z|u)p(u|y)}{p(z|y)}du\\ =-\int_{-\infty}^{\infty}\frac{(z-u)}{\sigma_{o}^{2}}p(u|z,y)du\\ =\frac{(-z+E[u|z,y])}{\sigma_{o}^{2}}\end{split}$ (100)
Therefore, we have that
$E[u|z,y]=z+\sigma_{o}^{2}\frac{d}{dz}(log(p(z|y))$ (101)
@@
Integrating backwards in time from $T$ to 0 obtains
@@
[…]
$dv-2t\frac{1+r(1+t^{2})}{r(1+t^{2})^{2}}v\text{d}t=-\frac{2}{r}\frac{t}{1+t^{2}}y\text{d}t+\sqrt{2t}\text{d}\overline{\beta}_{t}$ (111)
Next, by multiplying by an integrating factor we obtain
$\frac{e^{\frac{1}{r}\frac{1}{1+t^{2}}}}{1+t^{2}}dv-2t\frac{e^{\frac{1}{r}\frac{1}{1+t^{2}}}}{1+t^{2}}\frac{1+r(1+t^{2})}{r(1+t^{2})^{2}}v\text{d}t=-\frac{e^{\frac{1}{r}\frac{1}{1+t^{2}}}}{1+t^{2}}\frac{2}{r}\frac{t}{1+t^{2}}y\text{d}t+\sqrt{2t}\frac{e^{\frac{1}{r}\frac{1}{1+t^{2}}}}{1+t^{2}}\text{d}\overline{\beta}_{t}$ (112)
$d\Bigg{(}\frac{e^{\frac{1}{r}\frac{1}{1+t^{2}}}}{1+t^{2}}v\Bigg{)}=-\frac{e^{\frac{1}{r}\frac{1}{1+t^{2}}}}{1+t^{2}}\frac{2}{r}\frac{t}{1+t^{2}}y\text{d}t+\sqrt{2t}\frac{e^{\frac{1}{r}\frac{1}{1+t^{2}}}}{1+t^{2}}\text{d}\overline{\beta}_{t}$ (113)
Integrating backwards in time from $T$ to 0,
$\int_{T}^{0}d\Bigg{(}\frac{e^{\frac{1}{r}\frac{1}{1+t^{2}}}}{1+t^{2}}v\Bigg{)}=-\frac{2}{r}y\int_{T}^{0}e^{\frac{1}{r}\frac{1}{1+t^{2}}}\frac{t}{(1+t^{2})^{2}}\text{d}t+\int_{T}^{0}\sqrt{2t}\frac{e^{\frac{1}{r}\frac{1}{1+t^{2}}}}{1+t^{2}}\text{d}\overline{\beta}_{t}$ (114)
obtains the following
$e^{\frac{1}{r}}v(0)-\frac{e^{\frac{1}{r}\frac{1}{1+T^{2}}}}{1+T^{2}}v(T)=[e^{\frac{1}{r}}-e^{\frac{1}{r(1+T^{2})}}]y+\int_{T}^{0}\sqrt{2t}\frac{e^{\frac{1}{r}\frac{1}{1+t^{2}}}}{1+t^{2}}\text{d}\overline{\beta}_{t}$ (115)
We now solve for $v(0)$ to obtain
$v(0)=\frac{e^{-\frac{1}{r}\frac{T^{2}}{1+T^{2}}}}{1+T^{2}}v(T)+[1-e^{-\frac{T^{2}}{r(1+T^{2})}}]y+e^{-\frac{1}{r}}\int_{T}^{0}\sqrt{2t}\frac{e^{\frac{1}{r}\frac{1}{1+t^{2}}}}{1+t^{2}}\text{d}\overline{\beta}_{t}$ (116)
[…]
@@
Integrating backwards in time, from $T$ to 0,